OpenStudy (anonymous):

(x^2-100)÷(2x+10) use long division

OpenStudy (anonymous):

,,No it would be x-10. Check: (x+10)(x-10) Front=x^2 Outer=-10x Inner=10x Last=-100 Add them up and it equals x^2-100.